# Peeling Context from Cause for Multimodal Molecular Property Prediction

### v1

@@ 3.4 Case study
Figure 5: Layer-wise attribution maps in the multimodality setting across four benchmark datasets. Each row corresponds to a dataset and the molecular property of interest, indicated by the green arrow on the left. Three regimes are shown: under-peeling (left), optimal (center), and over-peeling (right). Atom-level contributions are highlighted with red atoms representing features most attributed to increasing the target property and blue atoms representing less contributing features. Optimal peeling yields clearer and more interpretable attribution compared to under- and over-peeling.
At the optimal depth, CLaP concentrates attribution on chemically coherent drivers per task.
In A. ESOL dataset, polar nitro groups and heteroatoms consistently dominate, reflecting their strong role in hydrogen bonding and aqueous solubility, especially in water.
In B. FreeSolv dataset, hydrocarbon segments together with the thio-linker carry the highest weights, consistently aligning with sulfur’s tendency to raise local hydrophobicity and, consequently, the molecule’s overall lipophilicity.
In C. Lipo dataset, the phenolic OH substituents and ring heteroatoms elevate hydrophilicity, while the extended aromatic surface is down-weighted, capturing the trade-off between polar handles and nonpolar bulk, consistent with established lipophilicity trends.
Finally, in D. CycPeptMPDB dataset, contiguous aromatic and aliphatic side chains form lipophilic patches that promote passive permeation, whereas backbone hydrogen-bond donors and acceptors contribute less, consistent with the desolvation cost of membrane crossing. This pattern reflects the intended peel—the shallow layers under-localize the signal, the deep layers over-peel and smear attribution, and the middle layers isolate chemically relevant substructures that drive the property.
4 Related Work
Graph causal inference.
Recent work aims to make graph predictors robust and interpretable by isolating label-relevant structure from spurious context. DIR selects invariant rationales (subgraphs that remain predictive across environments) and then predicts with the remainder (DIR). It depends on constructed environments from splits or augmentations, which can introduce artifacts not present in raw data. CIGA samples subgraphs with a learnable generator and enforces risk invariance with an information bottleneck (CIGA). The sampling loop can slow and destabilize training, and a bias toward small subgraphs risks missing key components. For molecules, environment modeling adds or replaces functional groups and treats these variants as environments (li2025soft), which may be chemically unfaithful and can widen the gap between training and test data even when raw data already mix causal and contextual factors. While CAL mitigates shortcuts and enhances interpretability (CAL), existing studies largely address classification. Regression is more challenging because the model must localize causal structure rather than exploit class-level artifacts. However, instead of crafting synthetic environments, our proposed CLaP achieves batch-wise invariance via the natural mini-batch reshuffling that occurs during training, yielding stable within-batch correlations across depths and better causal capture under re-batching.
Molecular property prediction.
Graph neural networks have become a cornerstone for molecular property prediction (Gilmer2017Neural; Wu2018MoleculeNet), achieving high accuracy on benchmarks by learning rich representations of molecular structures. Recent architectures extend beyond classical message passing: transformer-based models like the Molecule Attention Transformer incorporate geometric biases (inter-atomic distances) to capture 3D context (MAT), and multi-modal frameworks fuse complementary inputs (e.g., chemical language and graphs) to further boost performance (Rollins2024MolPROP). While these advances improve predictive power, they generally operate as black boxes and do not disentangle true causal factors from spurious correlations, limiting interpretability and their usefulness for guiding molecular design (JimenezLuna2020XAI).
5 Discussion
@@
We presented CLaP, a layerwise causal–context peeling framework that realizes batch-wise invariance, improving predictive accuracy and delivering atom-level, chemically consistent explanations.
Looking ahead, this framework can expand in both scope and mechanics. Specifically, it extends beyond scalar regression to classification by swapping scalar label supervision for within-class correlation objectives, while preserving depthwise alignment. Furthermore, it is not confined to molecular graphs and can be embedded in transformer backbones to route tokens into causal and trivial paths, yielding token level causal maps across modalities and domains. These extensions aim to better capture underlying causal structure and enable more faithful problem analysis.
Reproducibility statement
Implementation details and hyperparameters are in Appendix D.1. The supplementary material includes a code archive with configs, preprocessing, data loaders, architectures, and evaluations.
Ethics statement
We adhere to ICLR Code of Ethics. No human subjects or personally identifiable data are involved. To reduce misuse risk, we release benchmarking code and models only for these datasets and provide no synthesis or activity oracles. No conflicts of interest or sensitive sponsorships are present.
@@ Appendix A Theoretical Analysis
A.1 Batch-Wise Invariance Analysis
@@
$\mathrm{Corr}_{e}(s,y)$ $=\frac{\mathrm{Cov}_{e}(s,y)}{\sqrt{\mathrm{Var}_{e}(s)\,\mathrm{Var}_{e}(y)}}$ (13)
@@
$=\frac{a\,\sigma_{c,e}\,\sigma_{y,e}\,\kappa\;+\;b\,\sigma_{q,e}\,\sigma_{y,e}\,\alpha_{e}}{\sqrt{\,a^{2}\sigma_{c,e}^{2}+b^{2}\sigma_{q,e}^{2}+2ab\,\sigma_{c,e}\sigma_{q,e}\,\rho_{e}\,}\;\sigma_{y,e}}\quad\Big(\mathrm{Cov}_{e}(u,v)=\sigma_{u,e}\sigma_{v,e}\mathrm{Corr}_{e}(u,v)\Big)$
$=\frac{A_{e}\,\kappa+B_{e}\,\alpha_{e}}{\sqrt{A_{e}^{2}+B_{e}^{2}+2A_{e}B_{e}\rho_{e}}}\eqqcolon\frac{A_{e}\kappa+B_{e}\alpha_{e}}{D_{e}}.$ (14)
@@
$\frac{\partial}{\partial\alpha_{e}}\,\mathrm{Corr}_{e}(s,y)=\frac{\partial}{\partial\alpha_{e}}\,\frac{A_{e}\kappa+B_{e}\alpha_{e}}{D_{e}}=\frac{B_{e}}{D_{e}}.$ (15)
@@
$\left.\frac{\partial}{\partial B}\,\mathrm{Corr}_{e}(s,y)\right|_{B=0}=\Phi_{e}^{\prime}(0)=\frac{\alpha_{e}A_{e}^{2}-A_{e}\kappa(A_{e}\rho_{e})}{|A_{e}|^{3}}=\frac{\alpha_{e}-\kappa\rho_{e}}{|A_{e}|}.$ (16)
@@
$\left.\frac{\partial}{\partial b}\,\mathrm{Corr}_{e}(s,y)\right|_{b=0}=\sigma_{q,e}\,\left.\frac{\partial}{\partial B}\,\mathrm{Corr}_{e}(s,y)\right|_{B=0}=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\rho_{e}\big).$ (17)
@@
$\delta_{e}(b)=\delta_{e}(0)+b\,\Gamma_{e}+o(b),\qquad\Gamma_{e}=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\rho_{e}\big).$ (18)
@@
$b^{\star}=-\,\frac{\sum_{e}\delta_{e}(0)\,\Gamma_{e}}{\sum_{e}\Gamma_{e}^{2}}\;+\;o(1).$ (19)
@@
and since $\|\Gamma\|_{2}>0$ by (A2)–(A3), it follows that $b^{\star}\Rightarrow\ 0$ as the model gradually reduces the contribution of the trivial component. Hence in the limit the predictor discards the batch-dependent $q$ (i.e., $b^{\star}\Rightarrow\ 0$ so $s\approx\ a\,c$), so
@@
A.2 Trivial Branch Analysis
With squared loss and the stop–gradient in Eq. equation 11, training $t$ treats $y_{\mathrm{c}}^{\star}$ as a constant with respect to its parameters, so the trivial head solves the conditional least-squares problem
$t^{\star}(x)=\arg\min_{t}\ \mathbb{E}\!\big[(y-y_{\mathrm{c}}^{\star}-t(x))^{2}\mid x\big]=\mathbb{E}\!\big[y-y_{\mathrm{c}}^{\star}\mid x\big].$
@@ A.2 Trivial Branch Analysis
Let $r:=y-y_{\mathrm{c}}^{\star}$. By the definition of MSE risk $\mathcal{R}(f)=\mathbb{E}[(y-f(x))^{2}]$,
$\mathcal{R}(y_{\mathrm{c}}^{\star})=\mathbb{E}[r^{2}].$
@@
$\mathcal{R}(y_{\mathrm{c}}^{\star})=\mathbb{E}[r^{2}]=\mathbb{E}\!\big[\mathrm{Var}(r\mid x)\big]+\mathbb{E}\!\big[(\mathbb{E}[r\mid x])^{2}\big],$
which splits the error into (i) irreducible noise $\mathbb{E}[\mathrm{Var}(r\mid x)]$ and (ii) predictable bias $\mathbb{E}[(\mathbb{E}[r\mid x])^{2}]$.
Plugging $t^{\star}(x)=\mathbb{E}[r\mid x]$ into the predictor $\widehat{y}=y_{\mathrm{c}}^{\star}+t$ removes the predictable bias:
$\mathcal{R}(y_{\mathrm{c}}^{\star}+t^{\star})=\mathbb{E}\!\Big[\mathrm{Var}(r\mid x)+(\mathbb{E}[r\mid x]-t^{\star}(x))^{2}\Big]\overset{t^{\star}=\mathbb{E}[r\mid x]}{=}\mathbb{E}\!\big[\mathrm{Var}(r\mid x)\big].$
So adding $t^{\star}$ strictly reduces risk by $\mathbb{E}[(\mathbb{E}[r\mid x])^{2}]\!\geq\!0$ whenever the residual is partially predictable from $x$.
A.3 Target-Correlation Analysis
@@
$y=\theta\,c(x)+\eta,\qquad\mathbb{E}[\eta\mid x]=0,\quad\mathrm{Var}(\eta)>0,\quad\theta\geq 0$ (20)
@@
$\mathrm{Corr}(c,y)=\frac{\mathrm{Cov}(c,y)}{\sqrt{\mathrm{Var}(c)\mathrm{Var}(y)}}=\frac{\theta\,\mathrm{Var}(c)}{\sqrt{\mathrm{Var}(c)\,(\theta^{2}\mathrm{Var}(c)+\mathrm{Var}(\eta))}}=:\kappa\;<\;1.$ (21)
@@
$\sigma_{c,e}^{2}=\mathrm{Var}_{e}(c),\quad\sigma_{q,e}^{2}=\mathrm{Var}_{e}(q),\quad\kappa=\mathrm{Corr}_{e}(c,y)\ ,\quad\alpha_{e}=\operatorname{Corr}_{e}(q,y),\ \rho_{e}=\operatorname{Corr}_{e}(c,q),\ A_{e}=a\,\sigma_{c,e}$
@@
$\left.\frac{\partial}{\partial b}\operatorname{Corr}_{e}(s,y)\right|_{b=0}=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\,\rho_{e}\big)\ \eqqcolon\ \Gamma_{e},\qquad\left.\frac{\mathrm{d}}{\mathrm{d}b}\mathcal{L}_{\rho}\right|_{b=0}=2(\kappa-\rho_{\star})\sum_{e}\Gamma_{e}.$ (22)

### v2

@@ 3.4.2. Causal Invariance under Re-batching
Figure 5. Layer-wise attribution maps in a multimodal setting across four benchmark datasets. Each row corresponds to a dataset and target property (green arrow). Three regimes are shown: under-peeling (left), optimal (center), and over-peeling (right). Atom-level contributions are visualized, with red indicating positive attribution to the target property and blue indicating lower contribution. Optimal peeling yields clearer and more interpretable attribution than under- or over-peeling.
3.5. Case study
3.5.1. Layer-wise Causal Attribution
The causal split mechanism generates causal weights for each input component. Each causal block applies node-wise gating on atoms, producing $\alpha_{i}^{(\ell,m)}\!\in\![0,1]$ for atom/group $i$ in modality $m\!\in\!\{\textit{SM},\textit{PE},\textit{GE}\}$ at depth $\ell$. We take the final-layer weight as the causal score, $\pi_{i}^{(m)}=\alpha_{i}^{(L,m)}\in[0,1]$, and render causal saliency maps $\{\pi_{i}^{(m)}\}$ to inspect behavior.
Higher $\pi_{i}^{(m)}$ indicates atoms routed through the causal branch at the final layer. Figure 5 compares under-peeling, optimal-peeling, and over-peeling regimes. In the optimal regime, saliency concentrates on chemically meaningful moieties, consistent with chemical intuition. Across datasets, the resulting saliency patterns align with known structure–property relationships. A. ESOL emphasizes polar nitro groups and heteroatoms, reflecting their role in hydrogen bonding and aqueous solubility. B. FreeSolv assigns higher weight to hydrocarbon segments and the thio-linker, consistent with increased hydrophobicity. C. Lipo highlights phenolic OH substituents and ring heteroatoms while down-weighting extended aromatic surfaces, capturing the balance between polar functionalities and nonpolar bulk. D. CycPeptMPDB attributes importance to contiguous aromatic and aliphatic side chains that promote passive permeation, with reduced emphasis on backbone hydrogen-bond donors and acceptors. Overall, these trends illustrate the intended peeling behavior. Shallow layers under-localize attribution, deep layers over-peel and diffuse it, while intermediate layers isolate the chemically relevant substructures that drive each property.
3.5.2. Counterfactual Study
To further validate the learned causal weights, we perform a counterfactual experiment by substituting the highest-weight atoms with more soluble functional groups (Figure 6), and examine both solubility shifts and causal-weight responses. In the original molecule, the Cl atoms receive the highest causal scores, indicating that the model identifies them as solubility determinants. Replacing Cl with H removes these high-impact sites and leads to an increase in solubility. In contrast, substituting Cl with an O--H group preserves the same causal positions but reverses their chemical effect, producing an even larger increase. These counterfactual edits demonstrate that the model not only locates scaffold-level “causal sites,” but also correctly predicts both the direction and magnitude of property changes induced by specific chemical substitutions, reflecting counterfactual reasoning (Wang et al., 2024a).
Figure 6. Atom-level counterfactual explanation and attribution for ESOL solubility prediction produced by CLaP.
4. Related Work
4.1. Graph causal inference.
In graph causal inference, recent work aims to improve robustness and interpretability in graph learning by separating label-relevant causal structure from spurious contextual signals (An et al., 2025; Yao et al., 2024; Li et al., 2024). DIR selects invariant rationales—subgraphs that remain predictive across environments—and then predicts with the remainder (Wu et al., 2022). CIGA instead samples subgraphs via a learnable generator and enforces risk invariance through an information bottleneck (Chen et al., 2022).
In molecular settings, environment construction is simulated by adding or substituting functional groups and treating these variants as separate environments (Li et al., 2025b). However, manipulations may be chemically unrealistic and exacerbate the distribution gap between training and testing data. CAL alleviates shortcut learning and improves interpretability (Sui et al., 2022), yet prior work focuses on classification. Regression is inherently more challenging: the model must precisely localize causal substructures rather than rely on class-level artifacts or thresholds. Most recently, CGL extends causal graph learning to regression by jointly modeling causal and confounding subgraphs, with a particular focus on molecular property prediction (Yin et al., 2025).
However, instead of crafting synthetic environments, our proposed CLaP achieves batch-wise invariance via the natural mini-batch reshuffling during training, yielding stable within-batch correlations across depths and better causal capture under re-batching.
4.2. Molecular property prediction.
Neural architectures have become a cornerstone for molecular property prediction (Gilmer et al., 2017; Wu et al., 2018), achieving performance by learning molecular representations. Beyond classical message passing, recent architectures incorporate complex structural priors: transformer-based models such as MAT encode geometric biases via inter-atomic distances (Maziarka et al., 2019), while multimodal frameworks fuse chemical language with graph structure for richer cross-domain context (Rollins et al., 2024; Vinh et al., 2025b; Liu et al., 2024; Luo and Deng, 2025).
Despite these advances, most models behave as black boxes, capturing correlations without disentangling causal factors, which limits interpretability and usefulness in rational molecular design (Jiménez-Luna et al., 2020; Vinh et al., 2025a). To address this, several recent works enhance representation quality by incorporating structural insights or chemical priors. GSL-MPP leverages relational structure across molecules to improve predictive performance (Zhao et al., 2024), while MolFCL introduces fragment-aware contrastive learning and functional-group-level priors to obtain chemically robust and interpretable representations (Tang et al., 2025).
5. Discussion
@@
We introduced CLaP, a layerwise causal–context peeling framework that enforces batch-wise invariance to separate label-relevant structure from contextual shortcuts. This design improves predictive performance while yielding consistent atom-level explanations.
Beyond the current setting, the principle of depthwise causal alignment is broadly applicable. The framework can extend to classification via within-class correlation objectives and integrate into transformer architectures to route tokens into causal and context pathways, offering a general mechanism for interpretable and robust representation learning across structured prediction tasks.
Appendix Contents
1. Theoretical Analysis ........................................................................................................................................................................A.1
  1.1 Batch-Wise Invariance Analysis ........................................................................................................................................................................A.1
  1.2 Context Branch Analysis ........................................................................................................................................................................A.2
  1.3 Target-Correlation Analysis ........................................................................................................................................................................A.3
2. Additional Experiments ........................................................................................................................................................................B
  2.1 Effect of Peeling Hyperparameters ........................................................................................................................................................................B.1
  2.2 Additional Ablation Study ........................................................................................................................................................................B.2
3. Additional Case Studies and Multimodality Analysis ........................................................................................................................................................................7
  3.1 ESOL Dataset ........................................................................................................................................................................7
  3.2 FreeSolv Dataset ........................................................................................................................................................................9
  3.3 Lipo Dataset ........................................................................................................................................................................11
  3.4 CycPeptMPDB Dataset ........................................................................................................................................................................13
4. Implementation Details ........................................................................................................................................................................D.1
  4.1 Proposed Framework Implementation Details ........................................................................................................................................................................D.1
  4.2 Baselines Implementation Details ........................................................................................................................................................................D.2
@@ Appendix A Theoretical Analysis
A.1. Batch-Wise Invariance Analysis
@@
(10) $\mathrm{Corr}_{e}(s,y)$ $=\frac{\mathrm{Cov}_{e}(s,y)}{\sqrt{\mathrm{Var}_{e}(s)\,\mathrm{Var}_{e}(y)}}$
@@
(11) $=\frac{a\,\sigma_{c,e}\,\sigma_{y,e}\,\kappa\;+\;b\,\sigma_{q,e}\,\sigma_{y,e}\,\alpha_{e}}{\sqrt{\,a^{2}\sigma_{c,e}^{2}+b^{2}\sigma_{q,e}^{2}+2ab\,\sigma_{c,e}\sigma_{q,e}\,\rho_{e}\,}\;\sigma_{y,e}}$
(12) $=\frac{A_{e}\,\kappa+B_{e}\,\alpha_{e}}{\sqrt{A_{e}^{2}+B_{e}^{2}+2A_{e}B_{e}\rho_{e}}}\eqqcolon\frac{A_{e}\kappa+B_{e}\alpha_{e}}{D_{e}}.$
@@
(13) $\frac{\partial}{\partial\alpha_{e}}\,\mathrm{Corr}_{e}(s,y)=\frac{\partial}{\partial\alpha_{e}}\,\frac{A_{e}\kappa+B_{e}\alpha_{e}}{D_{e}}=\frac{B_{e}}{D_{e}}.$
@@
(14) $\left.\frac{\partial}{\partial B}\,\mathrm{Corr}_{e}(s,y)\right|_{B=0}=\Phi_{e}^{\prime}(0)=\frac{\alpha_{e}A_{e}^{2}-A_{e}\kappa(A_{e}\rho_{e})}{|A_{e}|^{3}}=\frac{\alpha_{e}-\kappa\rho_{e}}{|A_{e}|}.$
@@
(15) $\left.\frac{\partial}{\partial b}\,\mathrm{Corr}_{e}(s,y)\right|_{b=0}=\sigma_{q,e}\,\left.\frac{\partial}{\partial B}\,\mathrm{Corr}_{e}(s,y)\right|_{B=0}=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\rho_{e}\big).$
@@
(16) $\delta_{e}(b)=\delta_{e}(0)+b\,\Gamma_{e}+o(b),\qquad\Gamma_{e}=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\rho_{e}\big).$
@@
(17) $b^{\star}=-\,\frac{\sum_{e}\delta_{e}(0)\,\Gamma_{e}}{\sum_{e}\Gamma_{e}^{2}}\;+\;o(1).$
@@
and since $\|\Gamma\|_{2}>0$ by (A2)–(A3), it follows that $b^{\star}\Rightarrow\ 0$ as the model gradually reduces the contribution of the context component. Hence in the limit the predictor discards the batch-dependent $q$ (i.e., $b^{\star}\Rightarrow\ 0$ so $s\approx\ a\,c$), so
@@
A.2. Context Branch Analysis
With squared loss and the stop–gradient in Eq. equation 8, training $t$ treats $y_{\mathrm{c}}^{\star}$ as a constant with respect to its parameters, so the context head solves the conditional least-squares problem
$t^{\star}(x)=\arg\min_{t}\ \operatorname{\mathbb{E}}\!\big[(y-y_{\mathrm{c}}^{\star}-t(x))^{2}\mid x\big]=\operatorname{\mathbb{E}}\!\big[y-y_{\mathrm{c}}^{\star}\mid x\big].$
@@ A.2. Context Branch Analysis
Let $r:=y-y_{\mathrm{c}}^{\star}$. By the definition of MSE risk $\mathcal{R}(f)=\operatorname{\mathbb{E}}[(y-f(x))^{2}]$,
$\mathcal{R}(y_{\mathrm{c}}^{\star})=\operatorname{\mathbb{E}}[r^{2}].$
@@
$\mathcal{R}(y_{\mathrm{c}}^{\star})=\operatorname{\mathbb{E}}[r^{2}]=\operatorname{\mathbb{E}}\!\big[\operatorname{Var}(r\mid x)\big]+\operatorname{\mathbb{E}}\!\big[(\operatorname{\mathbb{E}}[r\mid x])^{2}\big],$
which splits the error into (i) irreducible noise $\operatorname{\mathbb{E}}[\operatorname{Var}(r\mid x)]$ and (ii) predictable bias $\operatorname{\mathbb{E}}[(\operatorname{\mathbb{E}}[r\mid x])^{2}]$.
Plugging $t^{\star}(x)=\operatorname{\mathbb{E}}[r\mid x]$ into the predictor $\widehat{y}=y_{\mathrm{c}}^{\star}+t$ removes the predictable bias:
$\mathcal{R}(y_{\mathrm{c}}^{\star}+t^{\star})=\operatorname{\mathbb{E}}\!\Big[\operatorname{Var}(r\mid x)+(\operatorname{\mathbb{E}}[r\mid x]-t^{\star}(x))^{2}\Big]\overset{t^{\star}=\operatorname{\mathbb{E}}[r\mid x]}{=}\operatorname{\mathbb{E}}\!\big[\operatorname{Var}(r\mid x)\big].$
So adding $t^{\star}$ strictly reduces risk by $\operatorname{\mathbb{E}}[(\operatorname{\mathbb{E}}[r\mid x])^{2}]\!\geq\!0$ whenever the residual is partially predictable from $x$.
A.3. Target-Correlation Analysis
@@
(18) $y=\theta\,c(x)+\eta,\qquad\operatorname{\mathbb{E}}[\eta\mid x]=0,\quad\operatorname{Var}(\eta)>0,\quad\theta\geq 0$
@@
(19) $\mathrm{Corr}(c,y)=\frac{\operatorname{Cov}(c,y)}{\sqrt{\operatorname{Var}(c)\operatorname{Var}(y)}}=\frac{\theta\,\operatorname{Var}(c)}{\sqrt{\operatorname{Var}(c)\,(\theta^{2}\operatorname{Var}(c)+\operatorname{Var}(\eta))}}=:\kappa\;<\;1.$
@@
$\sigma_{c,e}^{2}$ $=\operatorname{Var}_{e}(c),\quad\sigma_{q,e}^{2}=\operatorname{Var}_{e}(q),\quad\kappa=\operatorname{Corr}_{e}(c,y),$
$\alpha_{e}$ $=\operatorname{Corr}_{e}(q,y),\quad\rho_{e}=\operatorname{Corr}_{e}(c,q),\quad A_{e}=a\,\sigma_{c,e}$
@@
(20) $\left.\frac{\partial}{\partial b}\operatorname{Corr}_{e}(s,y)\right|_{b=0}$ $=\frac{\sigma_{q,e}}{|A_{e}|}\,\big(\alpha_{e}-\kappa\,\rho_{e}\big)\ \eqqcolon\ \Gamma_{e},$
$\left.\frac{\mathrm{d}}{\mathrm{d}b}\mathcal{L}_{\rho}\right|_{b=0}$ $=2(\kappa-\rho_{\star})\sum_{e}\Gamma_{e}.$